OpenStudy (anonymous):

The chess club has 9 members. They voted for president, vice president, and secretary of the club. How many different outcomes are possible?

OpenStudy (anonymous):

9P3= 504

OpenStudy (anonymous):

can a person be more than one?

OpenStudy (anonymous):

thought it was three ..guess not.

OpenStudy (anonymous):

9!/6!3!

OpenStudy (anonymous):

@myko, why?

OpenStudy (anonymous):

maybe i was right? 9/3 =3 ? 3 outcome?

OpenStudy (anonymous):

its a combinatorie number \[\left(\begin{matrix}9 \\ 3\end{matrix}\right)\]

OpenStudy (anonymous):

it depends on whether 1 person can have more than 1 job

OpenStudy (anonymous):

don't think so

OpenStudy (anonymous):

in that case 9!9!9!/3!